A 1.65 kg falcon catches a 0.295 kg dove from behind in midair. What is their velocity after impact if the falcon's velocity is initially 29.5 m/s and the dove's velocity is 4.95 m/s in the same direction?
Answer #1

using conservation of momentum

m1 v1 + m2 v2 = ( m1 + m2) v

1.65* 29.5 + 0.295* 4.95 = ( 1.65 + 0.295) v

v = 25.776 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck